# Galled Perfect Transfer Networks

### v2

@@ Abstract
With the goal of recovering the simplest possible scenarios in this model, we introduce galled perfect transfer networks, which are PTNs that are galled trees. Such networks are useful for characters that are incompatible in terms of tree-like evolution, but that do fit in an almost-tree scenario. We provide polynomial-time algorithms for two problems: deciding whether one can add transfer edges to a tree to transform it into a galled PTN, and deciding whether a set of characters are galled-compatible, that is, they can be explained by some galled PTN. We also analyze a real dataset comprising of a bacterial species trees and KEGG functions as characters, and derive several conclusions on the difficulty of explaining characters in a galled tree, which provide several directions for future research.
@@ 1 Introduction
When a perfect phylogeny does not exist for a set of characters, one may instead consider network-like structures to explain this diversity. To this end, Nakhleh et al. introduced Perfect Phylogenetic Networks (PPNs) in [nakhleh, nakhlehtesis]. These networks allow multi-state characters and require that, for each character $C$, the network contains a tree in which nodes in the same state are connected. This is a powerful model that is, unfortunately, difficult to work with, since even deciding whether a known network explains a set of characters is NP-hard. We note that in parallel to our work, this question was studied in the context of binary characters on galled trees [warnow2024statistically], which we discuss in more detail below.
@@
Importantly, PPNs were introduced as trees with additional transfer edges. This implies knowledge of the vertical versus horizontal transmissions, and that these networks belong to the class of tree-based networks [Pons2018, Fischer2020] (in fact, they are LGT networks, see next section). In [ptn], we introduced perfect transfer networks (PTNs), a specialization of PPNs where characters satisfy the same assumptions as perfect phylogenies: they are binary, have a unique origin, and are assumed to always be inherited vertically once acquired. An important difference of PTNs is that one cannot choose which subtree of the network can be used to explain a character, as the tree of vertical inheritance is fixed. Because of this, as we showed in [ptn], even PPNs on binary characters are different from PTNs. In the latter, a character may or may not be transferred horizontally to other species. This can be useful for transferrable characters that are difficult to revert, such as material acquired horizontally from mitochondria or chloroplasts resulting from endosymbiotic events [Zachar2020, anselmetti2021gene], or in the case of metabolites when HGT plays a role in the generation of new metabolic pathways in bacteria [Goyal2022]. From an algorithmic standpoint, an advantage of the more restricted PTN model is that the problems of deciding whether a network explains a set of characters, or whether a tree can be augmented with transfers to do so, become polynomial-time solvable.
@@
Thus, PTNs are a promising model for HGT inference from characters. However, PTNs have demonstrated a tendency to introduce an excessive number of transfer events. In [ptn], we provide an algorithm that shows that any tree can be made to explain any set of characters by adding transfers, although the resulting networks may be overly complicated and bloated with HGT edges. This raises the question of explaining a set of characters with the simplest possible network. One way would be to build a network or augment a tree with a minimum number of transfers, which is an unsolved complexity problem. Another direction is to impose structural conditions on the desired network.
@@
Our contribution. In this work, we focus on the second direction. We explore the evolutionary structure that many consider as the simplest beyond trees, namely galled trees. These consist of networks in which all underlying cycles are independent, and were first used in the context of hybridization and recombination [gusfield2004optimal]. When they fit the data, galled trees are desirable because of their parsimonious nature and ease of interpretation. They are also a popular graph-theoretical structure that serves as a first-step towards the development of more structurally complex networks [Huson2009, Cardona2020]. In our case, characters that can be explained by a galled tree can be thought of “not quite tree-like, but almost”.
@@
We present galled perfect transfer networks (galled PTNs) which are galled tree-based networks with unidirectional edges that explain a set of characters. We then provide polynomial-time algorithms for two problems. In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN. We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them. During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network. Althought most of our theoretical results are intuitive, the detailed proofs are sometimes involved. To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for the full details.
@@
Related work. Aside from PPNs, other models have been proposed to explain characters via networks. In recombination networks [gusfield2014recombinatorics], characters are explained by an ancestral recombination graph (ARG) in which hybridation nodes represent recombination events through crossovers. A fundamental difference is that such hybrids do not consider a donor/recipient relationship whereas in HGTs, it can be important to distinguish between the parental and lateral acquisition. As we showed in [ptn, Figure 3], PTNs and recombination graphs explain different sets of characters, even on galled trees with a single transfer/hybridation event, the main reason being that crossover events are different from transfer events. Nonetheless, it is worth mentioning that in [gusfield2004optimal], the author shows how to reconstruct a galled ancestral recombination graph from a set of $m$ characters and $n$ taxa, if possible, in time $O(nm+n^{3})$. Using this approach, the aforementioned work of Warnow et al. [warnow2024statistically] shows how a galled tree can be reconstructed (or not) from a set of characters in the PPN model, where characters are interpreted as bipartitions of the trees contained in the network, achieving the same time complexity. In Section 2, we show how this formulation differs from ours. In a similar vein, in [vanIersel2010, kelk2012] the authors study the question of reconstructing a network that displays a set of characters in the softwired sense, meaning that for each character, some tree contained in the network contains it as a clade (characters are called clusters therein). In particular, one can reconstruct in polynomial time a level-1 network that explains a set of characters, if one exists. As level-1 networks are closely related to galled trees, it is likely that the latter can be used in our setting. But, as also argued in [ptn, Figure 2], softwired characters can explain different sets of characters than PTNs. The main reason is that distinguishing between horizontal and vertical edges does not allow choosing which tree of the network should be used to explain a character.
@@ 2 Preliminaries
A phylogenetic network, or simply a network for short, is a directed acyclic graph $N=(V,E)$ with one node $\rho(N)$ of in-degree zero, called the root. A node of in-degree one and outdegree zero is a leaf, and $L(N)$ denotes the set of leaves of $N$. A node of in-degree and out-degree $1$ is a subdivision node, which we allow. For $W\subseteq V$, $N-W$ is the directed graph obtained after removing $W$ and incident edges. An underlying cycle of $N$ is a set of nodes and edges that form a cycle when ignoring the edge directions. A network $N$ is a galled tree if no two distinct underlying cycles of $N$ contain a common node (see Figure 1.(c)).
@@
A tree $T$ is a network with no underlying cycle. We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$ (we may drop the $T$ subscript if unambiguous). Note that $v$ is an ancestor and descendant of itself. Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e., if neither $u\preceq v$ nor $v\preceq u$. The parent of a node $v$ in $T$ is $p_{T}(v)$. For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants. We may write $L_{T}(v)$ as a shorthand for $L(T(v))$, or just $L(v)$ if $T$ is understood. The set $L(v)$ is called a clade of $T$.
@@
An LGT network (where LGT comes from Lateral Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$. The tree $\mathcal{T}_{N}$ is called the support tree of $N$. The edges in $E_{S}$ are called support edges and the edges in $E_{T}$ are called transfer edges. A node that is the endpoint of a transfer edge is called a transfer node. We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$. We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$. The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $N$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$. For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$. For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
@@ 2.1 Perfect transfer networks
Let $T$ be a tree on leafset $\mathcal{S}$, and let $\mathcal{C}$ be a set of characters. We say that $T$ is galled-completable for $\mathcal{C}$ if there exists a galled LGT network $N$ that explains $\mathcal{S}$ and whose base tree is $T$. We call such a galled PTN a galled-completion of $T$. An example of this problem is shown on Figure 1. For a set of characters $\mathcal{C}$ on taxa $\mathcal{S}$, we say that $\mathcal{C}$ is galled-compatible if there exists a galled tree on leafset $\mathcal{S}$ that explains $\mathcal{C}$. Our problems of interest are the following:
@@ 2.2 Properties of galled PTNs
We begin by stating properties of galled PTNs that will be useful throughout. Let $N$ be a network that explains a set of characters $\mathcal{C}$. Observe that if a leaf $l$ does not possess a character $C\in\mathcal{C}$, then by the “never lost once acquired” condition, the parent of $l$ in $\mathcal{T}_{N}$ cannot possess $C$ either (otherwise, the parent would be required to transfer $C$ to $v$). In fact, no ancestor of $l$ in $\mathcal{T}_{N}$ can possess $C$. This lets us deduce a subset $F_{C}$ of nodes that forbid $C$, as follows:
$F_{C}(N)=\{v\in V(N):\exists l\in L_{\mathcal{T}_{N}}(v)\mbox{ such that $l\notin C$}\}.$
@@
We also describe two useful generic properties of galled LGT networks. The first one says that a node $v$ cannot have two descending transfer nodes that each go outside of the $v$ subtree (otherwise, $v$ would be part of two distinct cycles).
@@
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ in $\mathcal{T}_{N}$ that are transfer nodes (with $v\in\{x,y\}$ being possible). Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
@@
If $v$ has two descendants $x,y$ with transfer neighbors that are external to the $v$ subtree of $\mathcal{T}_{N}$, then the two outgoing transfers inevitably create two distinct cycles that both contain $v$, contradicting the galled property. ∎
@@
We prove this by contraposition. Let $x^{\prime}$ and $y^{\prime}$ be nodes that share a transfer edge with $x$ and $y$, respectively. Suppose that neither $x^{\prime}$ nor $y^{\prime}$ are contained in $\mathcal{T}_{N}(v)$. Write $lca(a,b)$ for the lowest common ancestor of $a,b$ in $\mathcal{T}_{N}$. By our supposition, in $\mathcal{T}_{N}$ we have that $v$ is on the path from $lca(x,x^{\prime})$ to $x$, and from $lca(y,y^{\prime})$ to $y$. This implies the existence of two underlying cycles in $N$: there is a cycle formed by $lca(x,x^{\prime})-x^{\prime}-x-lca(x,x^{\prime})$ and the cycle formed by $lca(y,y^{\prime})-y^{\prime}-y-lca(y,y^{\prime})$ (where the edges between and $x^{\prime},x$, and between $y^{\prime},y$ are used, and the rest use edges of $\mathcal{T}_{N}$). These cycles are distinct since they use a different (unique) transfer edge. However, they intersect at $v$, and thus $N$ is not galled. ∎∎
The next property states that if some $x$ is able to reach a node $y$ that does not descend from it, then $x$ must of course achieve this through a transfer edge.
{lemmarep}
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network. Let $x,y\in V$ be two nodes that are not comparable in $\mathcal{T}_{N}$ and such that $x$ reaches $y$ in $N$. Then there exists an edge $(x^{\prime},y^{\prime})\in E_{T}$ such that $x^{\prime}\preceq_{\mathcal{T}_{N}}x$ and $y\preceq_{\mathcal{T}_{N}}y^{\prime}$.
{proofsketch}
Consider the first transfer edge $(x^{\prime},y^{\prime})$ on a path from $x$ to $y$ in $N$. Note that $x^{\prime}$ must descend from $x$ in $\mathcal{T}_{N}$. If $y$ is not a descendant of $y^{\prime}$, then from $y^{\prime}$ the path needs to borrow another transfer edge that goes out of the $y^{\prime}$ subtree of $\mathcal{T}_{N}$. Thus, the $y^{\prime}$ subtree has two transfers with external endpoints, contradicting Lemma 2.2. ∎
@@
Suppose for a contradiction that such $(x^{\prime},y^{\prime})$ does not exist in $N$. Since $x$ reaches $y$ in $N$, we know that there must exist at least another node $w\in V(N)$ incomparable to both $x$ and $y$ that serves as an intermediary node from $x$ to $y$. More specifically, there exists a transfer edge with an endpoint in $\mathcal{T}_{N}(x)$ and whose receiving end is $w$. Nevertheless, $y$ is not a descendant of $w$ (otherwise, $w=y^{\prime}$ would end the proof), which implies the existence of a transfer edge from some descendant of $w$ to a node incomparable with $w$. However this creates a situation where $\mathcal{T}_{N}(w)$ is a subtree with two transfer edges that have their endpoints outside of the subtree, which is forbidden by Lemma 2.2 ∎∎

### v3

@@ Abstract
With the goal of recovering the simplest possible scenarios in this model, we introduce galled perfect transfer networks, which are PTNs that are galled trees. Such networks are useful for binary characters that are incompatible in terms of tree-like evolution, but that do fit in an almost-tree scenario. We provide polynomial-time algorithms for two problems: deciding whether one can add transfer edges to a tree to transform it into a galled PTN, and deciding whether a set of characters are galled-compatible, that is, they can be explained by some galled PTN. We also analyze a real dataset comprising of a bacterial species trees and KEGG functions as characters, and derive several conclusions on the difficulty of explaining characters in a galled tree, which provide several directions for future research.
@@ 1 Introduction
When a perfect phylogeny does not exist for a set of characters, one may instead consider network-like structures to explain this diversity. To this end, Nakhleh et al. introduced Perfect Phylogenetic Networks (PPNs) in [nakhleh, nakhlehtesis]. These networks allow multi-state characters and require that, for each character $C$, the network displays a tree in which nodes in the same state are connected. That is, it is possible to remove all but one of the incoming edges of each reticulation, then label the internal nodes of the resulting tree with a state, such that every state forms a connected component. This is a powerful model that is, unfortunately, difficult to work with, since even deciding whether a known network explains a set of characters is NP-hard [nakhlehtesis]. We note that in parallel to our work, this question was studied in the context of binary characters on galled trees [warnow2024statistically], which we discuss in more detail below.
@@
Importantly, PPNs were introduced as trees with additional transfer edges. This implies knowledge of the vertical versus horizontal transmissions, and that these networks belong to the class of tree-based networks [Pons2018, Fischer2020] (in fact, they are LGT networks, see next section). In [ptn], we introduced perfect transfer networks (PTNs), a specialization of PPNs where characters satisfy the same assumptions as perfect phylogenies: they are binary, have a unique origin, and characters can never be lost in the vertical descendants once they are acquired. The latter two conditions are often called the no-homoplasy and no-loss conditions, respectively. An important difference of PTNs is that one cannot choose which subtree of the network can be used to explain a character, as the tree of vertical inheritance is fixed. This can be useful for transferrable characters that are difficult to revert, such as material acquired horizontally from mitochondria or chloroplasts resulting from endosymbiotic events [Zachar2020, anselmetti2021gene], or in the case of metabolites when HGT plays a role in the generation of new metabolic pathways in bacteria [Goyal2022]. From an algorithmic standpoint, an advantage of the more restricted PTN model is that the problems of deciding whether a network explains a set of characters, or whether a tree can be augmented with transfers to do so, become polynomial-time solvable.
@@
Thus, PTNs are a promising model for HGT inference from characters. However, PTNs have demonstrated a tendency to introduce an excessive number of transfer events. In [ptn], we provide an algorithm that shows that any tree can be made to explain any set of characters by adding transfers, although the resulting networks may be overly complicated and bloated with HGT edges. This raises the question of explaining a set of characters with the simplest possible network. One way would be to build a network or augment a tree with a minimum number of transfers, or impose structural conditions on the desired network. Notably, there are several positive results on the reconstructibility of networks with bounded level that explain characters in the softwired sense, i.e., each character can be explained by switching off all but one incoming edges of reticulations [vanIersel2010, kelk2012]. However, in PTNs, the no-loss condition implies that edges of vertical inheritance below a character’s origin cannot be switched off, and adapting the techniques developed in [vanIersel2010, kelk2012] to this additional restriction does not appear to be immediate (also see related works below).
@@
Our contribution. In this work, we explore the evolutionary structure that many consider as the simplest beyond trees, namely galled trees. These are also known as (binary) level-1 networks, and consist of networks in which all underlying cycles are independent, and were first used in the context of hybridization and recombination [gusfield2004optimal]. When they fit the data, galled trees are desirable because of their parsimonious nature and ease of interpretation. They are also a popular graph-theoretical structure that serves as a first-step towards the development of more structurally complex networks [Huson2009, Cardona2020]. In our case, characters that can be explained by a galled tree can be thought of “not quite tree-like, but almost”.
@@
We present galled perfect transfer networks (galled PTNs) which are galled tree-based networks with unidirectional edges that explain a set of characters. We then provide polynomial-time algorithms for two problems. In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN. We show that this verification can be done in polynomial time through the usage of an auxiliary structure. We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them. We show that it is possible to reconstruct a tree that can be augmented into a galled tree for a set of characters in polynomial time. During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network. Although most of our theoretical results are intuitive, the detailed proofs are sometimes involved. To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for the full details.
@@
Related work. Aside from PPNs, other models have been proposed to explain characters via networks. In recombination networks [gusfield2014recombinatorics], characters are explained by an ancestral recombination graph (ARG) in which hybridation nodes represent recombination events through crossovers. A fundamental difference is that such hybrids do not consider a donor/recipient relationship whereas in HGTs, it can be important to distinguish between the parental and lateral acquisition. As we showed in [ptn, Figure 3], PTNs and recombination graphs explain different sets of characters, even on galled trees with a single transfer/hybridation event, the main reason being that crossover events are different from transfer events. Nonetheless, it is worth mentioning that in [gusfield2004optimal], the author shows how to reconstruct a galled ancestral recombination graph from a set of $m$ characters and $n$ taxa, if possible, in time $O(nm+n^{3})$. Using this approach, the aforementioned work of Warnow et al. [warnow2024statistically] shows how a galled tree can be reconstructed (or not) from a set of characters in the PPN model, where characters are interpreted as bipartitions of the trees contained in the network, achieving the same time complexity. In Section 2, we show how this formulation differs from ours.
In a similar vein, in [vanIersel2010, kelk2012] the authors study the question of reconstructing a network that displays a set of characters in the softwired sense, meaning that for each character, some tree contained in the network contains it as a clade (characters are called clusters therein). It is known that for any fixed $k$, one can reconstruct in polynomial time a level-$k$ network that explains a set of characters, if one exists. In particular, level-1 networks are closely related to galled trees, so it is possible that these approaches can be used in our setting. But, as also argued in [ptn, Figure 2], softwired characters can explain more sets of characters than PTNs, mainly because of the no-loss condition.
@@ 2 Preliminaries
A phylogenetic network, or simply a network for short, is a directed acyclic graph $N=(V,E)$ with one node $\rho(N)$ of in-degree zero, called the root. We use $V(N)$ and $E(N)$ to denote the sets of nodes and edges of $N$, respectively. We say that a node $u\in V(N)$ reaches a node $v\in V(N)$ if there exists a directed path from $u$ to $v$ in $N$. A node of in-degree one and outdegree zero is a leaf, and $L(N)$ denotes the set of leaves of $N$. A node of in-degree and out-degree $1$ is a subdivision node, which we allow. For $W\subseteq V$, $N-W$ is the directed graph obtained after removing $W$ and incident edges. An underlying cycle of $N$ is a set of nodes and edges that form a cycle when ignoring the edge directions. A network $N$ is a galled tree if no two distinct underlying cycles of $N$ contain a common node (see Figure 1.(c)). Observe that two cycles may contain the same nodes but not the same edges, in which case they are considered distinct.
@@
A tree $T$ is a network with no underlying cycle. We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$ (we may drop the $T$ subscript if unambiguous). Note that $v$ is an ancestor and descendant of itself. Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e., if neither $u\preceq v$ nor $v\preceq u$. The parent of a non-root node $v$ in $T$ is $p_{T}(v)$. We shall only use the notions of $\preceq$, ancestors, descendants, and incomparable on trees, as they are not defined on networks. For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants. We may write $L_{T}(v)$ as a shorthand for $L(T(v))$, or just $L(v)$ if $T$ is understood. The set $L(v)$ is called a clade of $T$.
@@
An LGT network (where LGT comes from Lateral Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$. The tree $\mathcal{T}_{N}$ is called the support tree of $N$. The edges in $E_{S}$ are called support edges and the edges in $E_{T}$ are called transfer edges. A node that is the endpoint of a transfer edge is called a transfer node. We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$. We also assume that transfer nodes have exactly one child in $\mathcal{T}_{N}$. The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $N$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$. For $v\in V(N)\setminus\{\rho(N)\}$, we use the shorthand $p_{N}(v):=p_{\mathcal{T}_{N}}(v)$ to denote the parent of $v$ in the support tree of $N$. For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
Let us emphasize that in an LGT network, the partition $\{E_{S},E_{T}\}$ is specified. That is, there may be multiple ways of defining a tree within $N$ along with transfer edges, but $\{E_{S},E_{T}\}$ gives the unique desired way to do so [^2] Notation-wise, it may be more accurate to define an LGT network as a triple $(V,E,f)$ where $f:E\rightarrow\{support,transfer\}$ specifies the type of each edge, but we prefer to use the more convenient partition notation as in [Cardona_2015].. Therefore, the support tree $\mathcal{T}_{N}$ is defined unambiguously. This contrasts with so-called tree-based networks, where a partition into support and transfer edges is known to exist, but is not specified and may not be unique. Also note that galled trees always admit such a partition and are thus tree-based, but calling them galled LGT networks clarifies that the partition into support and transfer edges is given.
@@ 2.1 Perfect transfer networks
Let $T$ be a tree on leafset $\mathcal{S}$, and let $\mathcal{C}$ be a set of characters. We say that $T$ is galled-completable for $\mathcal{C}$ if there exists a galled LGT network $N$ that explains $\mathcal{S}$ and whose base tree is $T$. We call such a galled PTN a galled-completion of $T$. An example of this problem is shown on Figure 1. For a set of characters $\mathcal{C}$ on taxa $\mathcal{S}$, we say that $\mathcal{C}$ is galled-compatible if there exists a galled LGT network on leafset $\mathcal{S}$ that explains $\mathcal{C}$. Our problems of interest are the following:
@@ 2.2 Properties of galled PTNs
We begin by stating properties of galled PTNs that will be useful throughout. Let $N$ be an LGT network that explains a set of characters $\mathcal{C}$. Observe that if a leaf $s$ does not possess a character $C\in\mathcal{C}$, then by the “never lost once acquired” condition, the parent of $s$ in $\mathcal{T}_{N}$ cannot possess $C$ either (otherwise, the parent would be required to transmit $C$ vertically to $v$). In fact, no ancestor of $s$ in $\mathcal{T}_{N}$ can possess $C$. This lets us deduce a subset $F_{C}$ of nodes that forbid $C$, as follows:
$F_{C}(N)=\{v\in V(N):\exists s\in L_{\mathcal{T}_{N}}(v)\mbox{ such that $s\notin C$}\}.$
As an example, consider the network $N$ in Figure 1.(b). Here, $F_{C_{1}}(N)$ consists of every node from the root to the rightmost leaf (including the root and the leaf), because the character is not in that rightmost leaf. The set $F_{C_{2}}(N)$ contains all nodes on a path from the root to one of the two rightmost leaves, because those two do not have $C_{2}$ (so here, $F_{C_{2}}(N)$ includes one transfer node). Note that the same logic applies to subfigure (c).
@@
We also describe two useful generic properties of galled LGT networks. The first one says that a node $v$ cannot have two descending transfer nodes that each go outside of the descendants of $v$ in the support tree (otherwise, $v$ would be part of two distinct cycles).
@@
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ in $\mathcal{T}_{N}$ that are transfer nodes (with $v\in\{x,y\}$ being possible). Then one of the transfer edges of $N$ incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
@@
Suppose that $v$ has two distinct descendants $x,y$ in $\mathcal{T}_{N}$ that are both transfer nodes in $N$. Let $x^{\prime}$ (respectively $y^{\prime}$) be the other endpoint of the transfer edge of $N$ containing $x$ (resp. $y$), that is, $(x,x^{\prime})\in E_{T}$ or $(x^{\prime},x)\in E_{T}$ (same for $y$ and $y^{\prime}$). If both $x^{\prime}$ and $y^{\prime}$ are not descendants of $v$ in $\mathcal{T}_{N}$, then the two transfer edges containing $x$ and $y$ go outside of $\mathcal{T}_{N}(v)$ and create two cycles that both contain $v$, and these cycles are distinct because they use a different transfer edge, contradicting the galled property. ∎
@@
We prove this by contraposition. Let $x^{\prime}$ and $y^{\prime}$ be nodes that share a transfer edge with $x$ and $y$, respectively. Suppose that neither $x^{\prime}$ nor $y^{\prime}$ are contained in $\mathcal{T}_{N}(v)$. Write $lca(a,b)$ for the lowest common ancestor of $a,b$ in $\mathcal{T}_{N}$. By our supposition, in $\mathcal{T}_{N}$ we have that $v$ is on the path from $lca(x,x^{\prime})$ to $x$, and from $lca(y,y^{\prime})$ to $y$. This implies the existence of two underlying cycles in $N$: there is a cycle formed by $lca(x,x^{\prime})-x-x^{\prime}-lca(x,x^{\prime})$ and the cycle formed by $lca(y,y^{\prime})-y-y^{\prime}-lca(y,y^{\prime})$ (where here, the dashes represent paths; the edges between and $x,x^{\prime}$, and between $y,y^{\prime}$ are used, and the rest use edges of $\mathcal{T}_{N}$). These cycles are distinct since they use a different (unique) transfer edge. However, they intersect at $v$, and thus $N$ is not galled. ∎∎
The next property states that if some $x$ is able to reach a node $y$ in $N$ but $y$ is not a vertical descendant of $x$, then $x$ must reach $y$ through a path that uses exactly one transfer edge.
Lemma 2.
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network. Let $x,y\in V$ be two nodes that are incomparable in $\mathcal{T}_{N}$ and such that $x$ reaches $y$ in $N$. Then there exists a transfer edge $(x^{\prime},y^{\prime})\in E_{T}$ such that $x^{\prime}\preceq_{\mathcal{T}_{N}}x$ and $y\preceq_{\mathcal{T}_{N}}y^{\prime}$.
@@
Consider the first transfer edge $(x^{\prime},y^{\prime})$ on a path from $x$ to $y$ in $N$, which must exist. Note that $x^{\prime}$ must descend from $x$ in $\mathcal{T}_{N}$. If $y$ is not a descendant of $y^{\prime}$ in $\mathcal{T}_{N}$, then from $y^{\prime}$ the path needs to borrow another transfer edge that goes out of the $y^{\prime}$ subtree of $\mathcal{T}_{N}$. Thus $y^{\prime}$ has two distinct descendants in $\mathcal{T}_{N}(y^{\prime})$ that are transfer nodes with external endpoints, namely itself and the next transfer node on the path, contradicting Lemma 2.2. ∎∎